OpenStudy (anonymous):

Please help me!!! Formula is d = ut + 0.5at2. Solve for a

OpenStudy (anonymous):

Thanks so much guys for coming here for help :)

OpenStudy (anonymous):

':3

OpenStudy (ybarrap):

What is the initial velocity? What is the time? What is the acceleration?

OpenStudy (anonymous):

time =t, u = velocity and acceleration = a

OpenStudy (anonymous):

hold on

OpenStudy (ybarrap):

You need, t,u and a to solve for d.

OpenStudy (anonymous):

donuts.....? lol :3 sorry, idk.

OpenStudy (anonymous):

I just don't get it how to do it :(

OpenStudy (shaik0124):

d-ut/0.5t^2=a

OpenStudy (anonymous):

How did you place it front from back? I meant ut and you also put the a in the end

OpenStudy (shaik0124):

first subtract ut on both sides u will get d-ut=ut-ut+0.5at2 now d-ut=0.5at2 divide both sides by 0.5t2 u will get d-ut/0.5t2=a

OpenStudy (shaik0124):

a=d-ut/0.5at2

OpenStudy (anonymous):

thank you so much for help!!!!!! I understand it very well now. thanks alot !!! :)

OpenStudy (shaik0124):

its my pleasure